Find the perimeter in cm of the rectangle shown below.
1st side= 2y
2nd side= 2x+1
3rd side= x+2
4th side= x+y+2

Answer by swincher4391(1107)

2y = x+2
2x+1 = x+y+2
2x+1 = y+(x+2)
2x+1 = 3y
x+2 = 2y
2x+1 =3y
Multiply top by 2.

2x+4 = 4y
2x+1 = 3y
---------
3= y
x+2 = 2*3
x = 4
1st side: 6 cm
2nd side: 9 cm
3rd side: 6 cm
4th side: 9 cm
Perimeter: 6+9+6+9 = 30 cm